OpenStudy (anonymous):

Franco bought 1.5 pounds of deli cheese for . If he went back for another 2 pounds of the same cheese, how much could Franco expect to pay?

Parth (parthkohli):

1.5 pounds of deli cheese for?

OpenStudy (anonymous):

$ 6.75

Parth (parthkohli):

Okay, so you'd solve it by: \(\Large \color{purple}{\rightarrow {6.75 \over 1.5} * 2 }\) \(\Large \color{purple}{\rightarrow {675 \over 150} * 2 }\) \(\Large \color{purple}{\rightarrow {135 \over 75} * 2 }\) \(\Large \color{purple}{\rightarrow {27 \over 15} * 2 }\) \(\Large \color{purple}{\rightarrow {54 \over 15} }\) \(\Large \color{purple}{\rightarrow 3{4 \over 15} }\) Se, 3 4/15 is the answer.

OpenStudy (anonymous):

"Franco bought 1.5 pounds of deli cheese for . If he went back for another 2 pounds of the same cheese, how much could Franco expect to pay" You want to find out how much the cheese is per pound so: $ Price/1.5 = Price Per Pound. PPP x 2 = your answer. "The scale on a map indicates that 2 cm is equivalent to 15 miles. The distance Christine needs to travel is 5 cm on the map. How many miles does Christine need to travel" If 2cm = 15 miles the 1 cm = 15/2 or 7.5 miles. 5cm = 5 x 7.5 "The ratio of children to adults at a baseball game is 2 to 5. If there are 7,200 children at the game, which of the following could not be used to determine the number of adults" There are 2 children for every 5 adults or 5/2 = 2.5 adults per child. So divide 7200 by 2 then multiply by 2.5.